# Fermion-Boson Stars as Attractors in Fuzzy Dark Matter and Ideal Gas Dynamics

### v1

@@ Abstract
The evolution of the condensation, at least that of the FDM as indicated in Levkov et al. (2018); Chen et al. (2021), can be followed through the maximum of the density that reveals how the core accretes mass from the surroundings until saturation. In Fig. 3 we show the maximum of $\rho_{FDM}$ and $\rho$ as function of time, for the two values of $MR=0.1,0.2$ and the three values of $K=0.1,1,10$. Blue solid lines and orange dashed lines represent the results for $MR=0.1$ and $MR=0.2$. In all cases, the maximum of $\rho_{FDM}$ grows as typically without the IG starting at condensation time $\tau_{g}$ Levkov et al. (2018); Chen et al. (2021), in our case $\tau_{g}\sim 7$, and from then on the condensation of a core starts happening. On the other hand, the density of the IG prior to $\tau_{g}$ is erratic and in fact decreases, which indicates a stage of nearly uniform distribution, however, after $\tau_{g}$ the IG density grows and stabilizes, which indicates that after the FDM starts condensing and its dominating gravitational potential gets deeper, the IG starts accumulating around there and stabilizes. This is confirmed by the snapshots in Fig. 1.
@@
The condensation process results in an FDM-IG core that settles into a virialized configuration of the Schrödinger-Poisson-Euler (SPE) system as shown next. Fig. 4 shows the evolution of the various energies involved for the simulations with $MR=0.1$. For the FDM component these energies include the kinetic energy $K_{\text{FDM}}=-\frac{1}{2}\int_{D}\Psi^{*}\nabla^{2}\Psi\,d^{3}x$, the gravitational energy $W_{\text{FDM}}=\frac{1}{2}\int_{D}|\Psi|^{2}V\,d^{3}x$, the total energy $E_{\text{FDM}}=K_{\text{FDM}}+W_{\text{FDM}}$, and the virial factor $Q_{\text{FDM}}=2K_{\text{FDM}}+W_{\text{FDM}}$. For the gas component, the scalars are the kinetic energy $K_{\text{IG}}=\frac{1}{2}\int_{D}\rho|\vec{v}|^{2}\,d^{3}x$, the gravitational energy $W_{\text{IG}}=\frac{1}{2}\int_{D}\rho V\,d^{3}x$, the internal energy $U_{\text{IG}}=\int_{D}\rho e\,d^{3}x$, the total energy $E_{\text{IG}}=K_{\text{IG}}+W_{\text{IG}}+U_{\text{IG}}$, and the virial factor $Q_{\text{IG}}=2K_{\text{IG}}+W_{\text{IG}}+3U_{\text{IG}}$. Similar results are found for the simulations with $MR=0.2$.
@@
The plots illustrate two key aspects of the evolution. First, the virial factor $Q=Q_{\text{FDM}}+Q_{\text{IG}}\approx 0$ for times $t>\tau_{g}$ indicates that the system has approached a dynamically stable equilibrium. This near-zero value of $Q$ is a reliable indicator of virialization. As a result, the overall structure no longer undergoes significant changes, indicating that the system has settled around stable configuration.
@@
A second lead is the decrease of the IG kinetic energy toward zero, which indicates not only the approach toward a stationary state but also the approach toward hydrostatic equilibrium. In hydrostatic equilibrium, the inward gravitational force is balanced by the outward pressure forces, resulting in a configuration stable against both collapse and expansion. The small kinetic energy suggests minimal bulk motion within the gas, indicating that the gas distribution has settled into a nearly steady-state configuration. Consequently, this relaxed configuration of FDM-IG, formed through the condensation process, transitions to a structure that maintains equilibrium under its self-gravity and pressure forces, which are characteristics of a self-gravitating, hydrostatically balanced stationary Newtonian FBS Álvarez Rios and Guzmán (2023).
In Fig. 5 we compare the angularly averaged $\rho_{FDM}$ and $\rho$ after relaxation, with the densities of a ground state solution of a Newtonian FBS for one of our simulations with $MR=0.1$ at time $t=100$. An additional information of FBS is that they are constructed assuming a polytropic EoS, and in the case of Fig. 5, the FBS that fits the densities has a polytropic constant $\sim 103.5$. This type of fitting suggests not only that the FDM-IG system reaches radial-hydrostatic equilibrium, but also that entropy is nearly conserved, highlighting the robustness of these solutions.
Concerning the properties of the final configurations, the results can be summarized as follows. For a given $MR$, the central density and core radius of the FDM is independent of the initial polytropic constant $K$ of the IG within a few percent; on the other hand, for the IG we find that for the bigger $K$, the central density is smaller, indicating that the bigger the $K$, the less compact the IG distribution within the FDM core. This diversity of FDM and IG distributions are contained within wide range of FBS solutions Álvarez Rios and Guzmán (2023).
Conclusions. The condensation of the dominant FDM induces the stabilization of the IG, driving the system toward a virialized state. The system evolves into a dynamically stable equilibrium, characterized by a virial factor $Q\approx 0$ and a reduction in the kinetic energy of the gas, suggesting hydrostatic equilibrium. This indicates that the system stabilizes under its own self-gravity and pressure forces, transitioning from an initial blob formed from random initial conditions, into a stationary configuration of the SPE system. The results show the formation of a FBS from an initially random distribution of FDM and an IG with an 80% and 90% of FDM mass, and could be extended to study the formation and phenomenology of FBS in scenarios where the IG dominates.
In the context of FDM, if the IG is to be considered as a simple model of luminous matter, the attractor nature of FBS would imply that the distribution of the IG associated to luminous matter, should be common in galactic cores. In this sense, FBS as cores can impose limitations to the FDM model. One alternative can be the study of the coupling of oscillations between the two components, that has been found happen for FBS Álvarez Rios and Guzmán (2023).
The attractor character of FBS can be further explored by examining dynamical evolutions from smooth initial conditions, such as spherical collapse models. Moreover, non-ideal gas properties of the ISM can be studied by adding cooling and heating terms to Eq. (4). This will provide crucial improvements to FDM halo models used in precision constraints on allowed boson masses Vogt et al. (2023).
Thinking beyond the FDM-IG attractor cores, other scenarios offer more freedom to obtain less symmetric configurations of FDM+IG based on interaction between varios structures, for example spiral, disk and bullet-shaped contours of IG can be constructed Álvarez-Rios and Guzmán (2022), which may provide the sufficient variety of observed luminous structures beyond FBS cores but within the FDM phenomenology.
Acknowledgments
Iván Álvarez receives support from the CONAHCyT graduate scholarship program. This research is supported by grants CIC-UMSNH-4.9, Laboratorio Nacional de Cómputo de Alto Desempeño Grant No. 1-2024, CONAHCyT Ciencia de Frontera 2019 Grant No. Sinergias/304001.

### v2

@@ Appendix A Units
Appendix B Further diagnostics
We also analyze the effect of the initial velocity dispersion on the condensation process. Fig. 4 shows the time evolution of the maximum values of $\rho_{FDM}$ and $\rho$ for the case with $MR=1$ and $K=0.1$, varying the initial momentum dispersion with $\sigma=0.5$, $1$, and $2$. The results indicate that condensation also occurs within this range of dispersions, but the dynamics and efficiency of the process are affected. In particular, smaller values of $\sigma$ lead to a more pronounced growth of $\rho_{FDM}^{\text{max}}$, as the initial configuration is colder and more prone to gravitational collapse. In contrast, higher values of $\sigma$ delay the condensation and lead to a smaller central density $\rho_{FDM}^{\text{max}}$. The IG for this small $K$ distributes with a declining density while approaching a virialized state.
Figure 4: Maximum of $\rho_{FDM}$ (left) and $\rho$ (right) as function of time. Blue, orange and green lines indicate the cases with mass ratios $\sigma=0.5$, $1$, and 10, respectively.
Appendix C Virialization
Figure 5: Time evolution of the FDM and IG properties as function of time for the simulations with $MR=1$. These plots illustrate the relaxation process of the FDM-IG system, that leads to the formation of a Newtonian FBS. For the FDM component, we show the kinetic energy $K_{\text{FDM}}$, gravitational energy $W_{\text{FDM}}$, total energy $E_{\text{FDM}}$, and the virial scalar $Q_{\text{FDM}}$, all normalized by the initial absolute total energy $|E_{\text{FDM}}(0)|$. For the IG component, we display the kinetic energy $K_{\text{IG}}$, gravitational energy $W_{\text{IG}}$, internal energy $U_{\text{IG}}$, total energy $E_{\text{IG}}$, and the virial scalar $Q_{\text{IG}}$, each one normalized by the initial absolute total energy $|E_{\text{IG}}(0)|$. These energy diagnostics emphasize the stabilization of the system into a virialized configuration, with both FDM and IG components approaching stable energy values over time. Finally $Q_{FDM}\sim 0$ and $Q_{IG}\sim 0$ with time, which indicates that the two components evolve near a virialized state separately. Similar results are found for $MR=0.1$ and 10 in the SM Alvarez-Rios et al. (2025).
@@ Appendix C Virialization
The condensation process results in an FDM-IG core that settles into a virialized configuration of the SPE system as shown next. Fig. 5 shows the evolution of the various energies for the simulations with $MR=1$ and Figs. 4 of the SM for $MR=0.1,10$ in the SM. For the FDM component these energies include the kinetic energy $K_{\text{FDM}}=-\frac{1}{2}\int_{D}\Psi^{*}\nabla^{2}\Psi\,d^{3}x$, the gravitational energy $W_{\text{FDM}}=\frac{1}{2}\int_{D}|\Psi|^{2}V\,d^{3}x$, the total energy $E_{\text{FDM}}=K_{\text{FDM}}+W_{\text{FDM}}$, and the virial scalar $Q_{\text{FDM}}=2K_{\text{FDM}}+W_{\text{FDM}}$. For the gas component, the scalars are the kinetic energy $K_{\text{IG}}=\frac{1}{2}\int_{D}\rho|\vec{v}|^{2}\,d^{3}x$, the gravitational energy $W_{\text{IG}}=\frac{1}{2}\int_{D}\rho V\,d^{3}x$, the internal energy $U_{\text{IG}}=\int_{D}\rho e\,d^{3}x$, the total energy $E_{\text{IG}}=K_{\text{IG}}+W_{\text{IG}}+U_{\text{IG}}$, and the virial scalar $Q_{\text{IG}}=2K_{\text{IG}}+W_{\text{IG}}+3U_{\text{IG}}$. Similar results are found for the simulations with $MR=1$ and 10.
@@
The plots illustrate two key aspects of the evolution. First, the virial scalar $Q=Q_{\text{FDM}}+Q_{\text{IG}}\approx 0$ for times $t>\tau_{g}$ indicates that the system has approached a dynamically stable equilibrium. This near-zero value of $Q$ is an indicator of virialization. As a result, the overall structure no longer undergoes significant changes, indicating that the system has settled around stable configuration.
@@
Another evidence is the decrease of the IG kinetic energy toward zero, which indicates not only the approach toward a stationary state but also the approach toward hydrostatic equilibrium. In hydrostatic equilibrium, the inward gravitational force is balanced by the outward pressure forces, resulting in a configuration stable against both collapse and expansion. The small kinetic energy suggests minimal bulk motion of the gas, indicating that the gas distribution has settled into a nearly steady-state configuration. Consequently, this relaxed configuration of FDM-IG, formed through the condensation process, towards a hydrostatically balanced FDM-IG system that now we compare with FBS solutions.